An acetic acid solution (Ka= 1.8x10^-5) with an unknown concentration has a pH of 2.98 what is the concentration of the acidic acid solution?
Answer #1

pH = 2.98

or

-log [H+] = 2.98

or

[H+] = 1.047 * 10^-3 M

thus

[CH3COO-] = 1.047 * 10^-3 M

let concentration of acid = x M

thus

Ka = [H+][CH3COO-] / [CH3COOH]

or

1.8 x 10^-5 = 1.047 * 10^-3 * 1.047 * 10^-3 / (x - 1.047 * 10^-3)

or

x = 0.062 M

thus concentration of acid = 0.062 M